OpenStudy (anonymous):

Find the coefficient of x^10 in the power series (x^2 +x^4+x^6+....)(x^3+x^6+x^9+.....)(x^4+x^8+x^12+....).Can someone help me?

OpenStudy (anonymous):

the ans is 0 since no solution for 2a+3b+4c=10, for any natural number a, b, c

OpenStudy (raden):

(x^2 +x^4+x^6+....)(x^3+x^6+x^9+.....)(x^4+x^8+x^12+....) = x^2(1+x^2+x^4+...)x^3(1+x^3+x^6+...)x^4(1+x^4+x^8+...) = x^9 (1+x^2+x^4+...)(1+x^3+x^6+...)(1+x^4+x^8+...) the multiply of terms in brackets cant getting the x term, so there is not a x^10, in other words @enka is right :)

OpenStudy (anonymous):

I don't understand. how can you lead to that equation? and even though if so, you still have so many cases to get that for example 6 [3*2] +0{3*0] +4 [4*1]

OpenStudy (anonymous):

ok . I got it now thanks alot. what if every terms has 1+.... in the front of the sequence?

OpenStudy (anonymous):

then, you'll get the ans

OpenStudy (anonymous):

I rewrite the whole thing (1+x^2+x^4+x^6+.....)(1+x^4+x^8+x^12+....)(1+x^6+x^12+x^18+....)

OpenStudy (anonymous):

the question is the same, coefficient of x^10

OpenStudy (tkhunny):

These can be a little tedious, but you can simplify a bit, depending on the exct question. It is sufficient to track this: (x^2 +x^4+x^6+x^8+x^10)(x^3+x^6+x^9)(x^4+x^8) Everything else will be over 10 Multiply the first two: x^5 + x^7 + x^8 + x ^9 + x^10 + a bunch of other stuff that we can discard. (x^5 + x^7 + x^8 + x ^9 + x^10)(x^4+x^8) Recognizing that the lowest exponent we brought with us in the last factor is 4, we can discard from the larger factor all but the first term (x^5)(x^4+x^8) x^9 + x^13 -- There is no x^10

OpenStudy (anonymous):

now the problem count back 3 post from this message

OpenStudy (tkhunny):

Better: It is sufficient to track this: (x^2 +x^4+x^6+x^8+x^10)(x^3+x^6+x^9)(x^4+x^8) Everything else will be over 10 Factor (x^9)(1 + x^2 + x^4 + x^6 + x^8)(1 + x^3 + x^6)(1 + x^4) And the question becomes, what is the coefficient of x^1 in the following expression. (1 + x^2 + x^4 + x^6 + x^8)(1 + x^3 + x^6)(1 + x^4) It should be relatively obvious that the desired coefficient is zero (0).

OpenStudy (anonymous):

It is not that. they are the infinite polynomial . they don't stop at any where

OpenStudy (anonymous):

the first ( ) =\[\frac{ 1 }{ 1-x^2}\]

OpenStudy (anonymous):

the second( ) =\[\frac{ 1 }{1-x^4 }\]

OpenStudy (anonymous):

the third ( ) =

OpenStudy (anonymous):

\[\frac{ 1 }{ 1-x^6 }\]

OpenStudy (tkhunny):

Okay, then (1 + x^2 + ...) *(1 + x^3 + ...) *(1 + x^4 + ...) Again, it is sufficient to track these: (1 + x^2 + x^4 + x^6 + x^8 + x^10) *(1 + x^3 + x^6 + x^9) *(1 + x^4 + x^8) Multiply the 1st 2 (1 + x^2 + x^3 + x^4 + x^5 + 2x^6 + x^7 + 2x^8 + 2x^9 + 2x^10 + other stuff that we can ignore) (1 + x^2 + x^3 + x^4 + x^5 + 2x^6 + x^7 + 2x^8 + 2x^9 + 2x^10) Piecemeal the last factor 1 * 2x^10 = 2x^10 x^4 * 2x^6 = 2x^10 x^8 * x^2 = x^10 and we get 5x^10 Like I said, these can be a little tedious.

OpenStudy (anonymous):

yes you are right. Thanks a lot. I have to copy it to figure out the new way to have the answer

OpenStudy (anonymous):

I am sorry when I cannot give you the medal for your answer since it is not an independent question and from the original problem I gave out the all of them . Terribly sorry and thanks a looooot

OpenStudy (tkhunny):

Infinite though they are, if you ask me about the coefficient of x^10, there is quite a bit of the infinite polynomial that will not have any impact and it can be simply ignored.

OpenStudy (anonymous):

now I learn a new way to solve the problem

OpenStudy (anonymous):

I will post another question in Algebra. are you interesting in?